OpenStudy (anonymous):

HOW DO YOU SOLVE THIS? ∛2 + 2∛128

OpenStudy (ajprincess):

\[\sqrt[3]{128}=\sqrt[3]{4^3*2}\] \[=4\sqrt[3]{2}\]

OpenStudy (ajprincess):

\(64=4*4*4=4^3\)

OpenStudy (ajprincess):

\[\sqrt[3]{2}+2\sqrt[3]{128}=\sqrt[3]{2}+2*4\sqrt[3]{2}=\sqrt[3]{2}+8\sqrt[3]{2}=?\]